Calculate the magnitude and direction of the electric field at a point P which is 30 cm to the right of a point charge Q = -3.0 x 10-6 C


1
Expert's answer
2022-02-13T12:15:32-0500

The electric field due the point charge

"E=k\frac{q}{r^2}=9*10^9*\frac{3.0*10^{-6}}{(0.3)^2}=3*10^5\:\rm V/m"

Direction is toward the charge.